a) Determine the missing frequencies of the following distribution given that the median


is 33.5 and themode is 34.0.


Class limits 0-9 10-19 20-29 30-39 40-49 50-59 60-69 Total


Freq. 4 16 f3 f4 f5 6 4 230


b) Compute the arithmetic mean.


c) Compute the value below which 25% of the observations lie.


d) Compute the value above which 25% of the observations lie

1
Expert's answer
2022-03-28T16:48:26-0400

a) missing frequencies




Cumulative Frequency total = 30+f3+f4+f5=230

f3+f4+f5=200...........(i)

median class is 29.5 - 39.5

"33.5=29.5+[\frac{\frac{230}{2}-(20+f3)}{f4}]*10"

"4=[\frac{115-20-f3}{f4}]*10"

"0.4=\frac{95-f3}{f4}"

"0.4f4+f3=95" ..........(ii)

modal class is 29.5 - 39.5

"34=29.5+{\frac{f4-f3}{(f4-f3)+(f4-f5)}}*10"

"4.5=[\frac{f4-f3}{f4-f3+f4-f5}]*10"

"0.45=\frac{f4-f3}{2f4-f3-f5}"

"0.45(2f4-f3-f5)=f4-f3"

"-0.1f4+0.55f3-0.45f5=0" .....................(iii)

Solving equations (i), (ii) and (iii) yields

"f3=55"

"f4=100"

"f5=45"


b) Arithmetic mean

"=\frac{(4*4.75)+(16*14.5)+(55*24.5)+(100*34.5)+45*44.5)+(6*54.5)+(4*64.5)}{230}=\frac{7636}{230}=33.2"


c) value below 25%

"\frac{25}{100}*230=57.5"


d) value above 25%

"100-57.5=42.5"